OpenStudy (anonymous):

What are the x-intercept(s) of the graph of y + 6 = x2 − x?

OpenStudy (anonymous):

For finding x intercept put y = 0: \[0 + 6 = x^2 - x\] \[\implies x^2 - x - 6 = 0\] Factorize it and find x..